OpenStudy (anonymous):

find a line for an equation passing through (1,2) and parallel to it with the line given:y=3x-2 ....and also perpendicular

OpenStudy (anonymous):

The equation of the line through (1,2) must use the slope from the other give equation. Since I am looking for a PARALLEL line the slopes will be the SAME. Hence m = 3 Now plug (1,2) and m=3 into the POINT-SLOPE formula and simplify So y-y1 =m(x-x1) y-2=3(x-1) y-2 =3x-3 y = 3x -1

OpenStudy (anonymous):

ok i get it thank you...but what about the perpendicular?

OpenStudy (anonymous):

perpendicular lines have OPPOSITE RECIPROCAL slopes so change the sign and flip that junt

OpenStudy (anonymous):

lol